# You Only Sample Once: Taming One-Step Text-To-Image Synthesis by Self-Cooperative Diffusion GANs

### v2

@@ 6.2 Ablation Studies
Effect of Consistency Loss and LPIPS Loss. We evaluate the effect of consistency loss and LPIPS loss in YOSO. We find that the both consistency loss and LPIPS loss can improve the image quality, as indicated by FID. But we notice that if removing consistency loss or LPIPS loss, the generation performance would not decrease significantly. This indicates the effectiveness of our proposed adversarial formulation.
Adversarial Divergence . The key design of our proposed YOSO is smoothing the distribution by using self-generated data instead of real data as ground truth to perform adversarial divergence. We evaluate the variant of using as adversarial divergence. As shown in Table 1, the performance of the variant is significantly worse than YOSO. Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant. As shown in Figure 4, the discriminator loss of the variant is not stable and is much smaller than ours. This indicates that our approach can effectively smooth the distribution to make the distributions of fake and real samples closer to each other, making discrimination harder and hence resulting in higher discriminator loss. Furthermore, we observe mode collapse in the training of the variant. We note that a similar case also happened in SDXL-Turbo, a strong one-step text-to-image model, which also performs adversarial divergence directly against real data (See Figure 5). These results indicate the superiority of the proposed approach.
Comparison to UFOGen. Similar to us, UFOGen does not directly perform adversarial divergence against real data either. However, they perform adversarial divergence over corrupted data (i.e., ). We argue that the formulation is less effective to construct one-step generation model since their adversarial divergence only indirectly matches the one-step generation at clean data. We implement the UFOGen on CIFAR-10 with our own. Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO. As shown in Table 1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID. This further demonstrates the superiority of our proposed approach. Through comparison with different adversarial divergence formulations, we find that our formulated adversarial divergence not only smooths the distribution to stabilize the training but also provides effective learning on one-step generation.
@@ 6.3 Text-to-image generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt- which is a diffusion transformer with 0.6B parameters. We note that 0.6B parameters can be demonstrated as lightweight among text-to-image generation models. For the GAN discriminator, we construct the latent discriminator by pre-trained Stable Diffusion 1.5 using the proposed construction introduced in Section 5.1. We switch the pre-trained PixArt- to v-prediction by the proposed technique introduced in Section 5.1, followed by training on the JourneyDB dataset [36] with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only around 10 A800 days to be trained.
@@
Evaluation. For image quality, we employ Aesthetic Score (AeS) [45] to evaluate. We also adopt the Human Preference Score v2.1 (HPS v2.1) [51] to evaluate the image-text alignment and human preference. The AeS is computed by an aesthetic score predictor trained on LAION [45] datasets, without considering the image-text alignment. HPS is trained to predict human preference given image-text pairs, considering the image-text alignment and human aesthetic. We mainly compare our model against the open-source state-of-the-art models, i.e., SD-Turbo [44] and SDXL-Turbo [44].
@@
Quantitative Results. The quantitative results are presented in Table 3. To provide a direct comparison with SD-Turbo, we also fine-tune a model based on pre-trained SD 1.5 on JourneyDB with the introduced technique. We would like to highlight that it is better to conduct training on LAION dataset [45] which is the training set used by SD 1.5. However, it is not available when we conducted this research. Fine-tuning SD 1.5 on other datasets suffers from data shift, facing a risk of catastrophic forgetting. As shown in Table 3, both YOSO-SD and YOSO-PixArt- clearly beat the SD-Turbo which is based on SD 2.1, a better SD compared to SD 1.5. It is also observed that YOSO-PixArt- is able to compete with SDXL-Turbo which is a much bigger model with 2.6B parameters.
@@
Qualitative Comparison. To provide more comprehensive insight in understanding the performance of our YOSO, we further provide the qualitative comparison in Figure 5. The comparison shows that: YOSO clearly beats SD-Turbo in terms of both image quality and prompt alignment. Specifically, the samples by SD-Turbo are blurry to some extent. Although SDXL-Turbo with a larger model size improves the sample quality compared to SD-Turbo, we find that the samples by YOSO is sharper than SDXL-Turbo to some extent, and we do better in prompt alignment. For example, in the first row of Figure 5 related to a Pikachu, both SD-Turbo and SDXL-Turbo cannot follow the prompt to generate angry expressions and red eyes, while YOSO follows the prompt more precisely. Moreover, it is important to highlight that SDXL-Turbo seems to have a serious issue in mode collapse. Specifically, the first row and third row of Figure 5 show that samples by SDXL-Turbo are really close to each other. This indicates that SDXL-Turbo performs poorly in mode cover. Overall, based on quantitative and qualitative results, we conclude that our YOSO is better in sample quality, prompt alignment, and mode cover compared with the state-of-the-art one-step text-to-image models (i.e., SD-Turbo and SDXL-Turbo).
@@ 6.3.1 Zero-Shot One-Step 1024 Resolution Generation
where and means the model weight. We let . As shown in Figure 1(b), the YOSO constructed in this way can generate high-quality images with 1024 resolution. Note that YOSO even changes the predicted objective to v-prediction. The impressive performance indicates the robust generalization ability of our YOSO.
@@ 6.3.2 Fine-tuning Stable Diffusion via LoRA
We provide experiments of full fine-tuning PixArt- and SD 1.5 on JourneyDB above. It is worth mentioning that JourneyDB is not in the training set of SD 1.5, and training on it will cause a distribution shift to some extent, which hinders the wide application in downstream tasks. Hence, we further conduct experiments on fine-tuning SD 1.5 via LoRA. As we perform LoRA fine-tuning, it is expected that it will be worse than full fine-tuning, we propose to use SDXL-Turbo to generate data for LoRA fine-tuning, which avoids the risk of data shift to some extent and is economically efficient. We call the model fine-tuned in this way as YOSO-LoRA. We present the quantitative results in Table 4 and qualitative results in Figure 6. It can be observed that YOSO-LoRA archives even better results than LCM with 4 steps and is even comparable to InstaFLow with full fine-tuning. And from Figure 6, YOSO-LoRA has a clearly better sample quality than LCM with 2 steps and InstaFlow. These results indicate that our algorithm enables performing one-step text-to-image synthesis with reasonable quality by LoRA fine-tuning. However, we still observe a little bit of distribution shift compared to stable diffusion in YOSO-LoRA. This issue can only be addressed by training on the LAION dataset [45] or generating training data by SD 1.5 which is expensive.
@@ 6.4 Application
One promising and attractive property of text-to-image DMs is that they can be used in multiple downstream tasks as the base model. In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of text-to-image DMs. Since the relevant applications are mostly developed for stable diffusion, we mainly show the results obtained by YOSO-LoRA in this section.
Image-to-image. As shown in Figure 6, our YOSO-LoRA is capable of performing high-quality image-to-image editing [33] in one step.
Compatibility with ControlNet and Different Base Models. As shown in Figure 6, our YOSO-LoRA is compatible with ControlNet [61], following the condition well. Our YOSO-LoRA is also compatible with different base models (e.g., dreamshaper and realistic vision) fine-tuned from SD 1.5, preserving their style well.
We find that when applying YOSO-LoRA to new base models, it fails to produce reasonable samples in one step. However, YOSO-SD succeeds in doing so when combined with different base models using Equation 8. Unfortunately, the style of the new base models cannot be preserved, as YOSO-SD also alters the style of the original SD 1.5 due to distribution shift in our training set.

### v3

@@ 6.2 Ablation Studies
Effect of Consistency Loss and LPIPS Loss. We evaluate the effect of consistency loss and LPIPS loss in YOSO. We find that both consistency loss and LPIPS loss can improve the image quality, as indicated by FID (see Tab. 1). However, we notice that if removing consistency loss or LPIPS loss, the generation performance would not decrease significantly. This indicates the effectiveness of our proposed adversarial formulation.
Adversarial Divergence . The key design of our proposed YOSO is smoothing the distribution by using self-generated data instead of real data as ground truth to perform adversarial divergence. We evaluate the variant of using as adversarial divergence. As shown in Tab. 1, the performance of the variant is significantly worse than YOSO. Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant. As shown in Fig. 4, the discriminator loss of the variant is not stable and is much smaller than ours. This indicates that our approach can effectively smooth the distribution to make the distributions of fake and real samples closer to each other, making discrimination harder and hence resulting in higher discriminator loss. Furthermore, we observe mode collapse in the training of the variant. We note that a similar case also happened in SDXL-Turbo, a strong one-step text-to-image model, which also performs adversarial divergence directly against real data (See Fig. 5). These results indicate the superiority of the proposed YOSO.
Comparison to UFOGen. Similar to us, UFOGen does not directly perform adversarial divergence against real data either. However, they perform adversarial divergence over corrupted data (i.e., ). We argue that the formulation is less effective in constructing one-step generation model since their adversarial divergence only indirectly matches the one-step generation at clean data. We implement the UFOGen on CIFAR-10 with our own. Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO. As shown in Tab. 1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID. This further demonstrates the superiority of our proposed YOSO. Through comparison with different adversarial divergence formulations, we find that our formulated adversarial divergence not only smooths the distribution to stabilize the training but also provides effective learning on one-step generation.
@@ 6.3 Text-to-Image Generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt- which is a diffusion transformer with 0.6B parameters. We note that 0.6B parameters can be demonstrated as lightweight among text-to-image DMs. For the GAN discriminator, we construct the latent discriminator by pre-trained SD 1.5 using the proposed construction introduced in Sec. 5.1. We switch the pre-trained PixArt- to v-prediction by the proposed technique introduced in Sec. 5.1, followed by training on the JourneyDB dataset [35] with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only 30k iterations and around 10 A800 days to be trained.
@@
Evaluation. For image quality, we employ Aesthetic Score (AeS) [45] to evaluate. We also adopt the Human Preference Score (HPS) v2.1 [53] to evaluate the image-text alignment and human preference. The AeS is computed by an aesthetic score predictor trained on LAION [45] datasets, without considering the image-text alignment. HPS is trained to predict human preference given image-text pairs, considering the image-text alignment and human aesthetic. We mainly compare our model against the open-source state-of-the-art (SOTA) models, i.e., SD-Turbo [44] and SDXL-Turbo [44].
@@
Quantitative Results. The quantitative results are presented in Tab. 3. To provide a direct comparison with SD-Turbo, we also fine-tune a model based on pre-trained SD 1.5 on JourneyDB with the introduced technique. We would like to highlight that it is better to conduct training on LAION dataset [45] which is the training set used by SD 1.5. However, it is not available when we conducted this research. Fine-tuning SD 1.5 on other datasets suffers from distribution shift, facing a risk of catastrophic forgetting. As shown in Tab. 3, both YOSO-SD and YOSO-PixArt- clearly beat the SD-Turbo which is based on SD 2.1, a better SD compared to SD 1.5. It is also observed that YOSO-PixArt- outperforms SDXL-Turbo which is a much larger model with 2.6B parameters.
@@
Qualitative Comparison. To provide more comprehensive insight in understanding the performance of our YOSO, we further provide the qualitative comparison in Fig. 5. The comparison shows that: YOSO clearly beats SD-Turbo in terms of both image quality and prompt alignment. Specifically, the samples by SD-Turbo are blurry to some extent. Although SDXL-Turbo with a larger model size improves the sample quality compared to SD-Turbo, we find that the samples by YOSO are sharper than SDXL-Turbo to some extent, and better in prompt alignment. For example, in the first row of Fig. 5 related to a Pikachu, both SD-Turbo and SDXL-Turbo cannot follow the prompt to generate angry expressions and red eyes, while YOSO follows the prompt more precisely. Moreover, it is important to highlight that SDXL-Turbo seems to have a serious issue in mode collapse. Specifically, the third row of Fig. 5 shows that samples by SDXL-Turbo are really close to each other. Overall, based on quantitative and qualitative results, we conclude that our YOSO is better in sample quality, prompt alignment, and mode cover compared with the SOTA one-step text-to-image models.
@@ 6.3.1 Zero-Shot One-Step 1024 Resolution Generation
where and means the model weight. We let in our experiments. As shown in Fig. 1(b), the YOSO constructed in this way can generate high-quality images with 1024 resolution. Note that YOSO even changes the predicted objective to v-prediction. The impressive performance indicates the robust generalization ability of our proposed YOSO.
@@ 6.3.2 Fine-tuning Stable Diffusion via LoRA
We further conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of proposed YOSO. As we perform LoRA fine-tuning, it is expected that it will be worse than full fine-tuning, we propose to use SDXL-Turbo to generate data for LoRA fine-tuning, which avoids the risk of distribution shift to some extent and is economically efficient. We call the model fine-tuned in this way as YOSO-LoRA. We only generate one image for each caption, avoiding the risk of mode collapse brought by SDXL-Turbo. We train YOSO-LoRA with only 64 batch size and 4000 iterations. We present the quantitative results in Tab. 4 and qualitative results in Fig. 6. It can be observed that YOSO-LoRA achieves even better results than LCM with 4 steps and is comparable to InstaFLow with full fine-tuning. And from Fig. 6, YOSO-LoRA has a clearly better sample quality than LCM with 2 steps and InstaFlow. These results indicate that our algorithm enables one-step text-to-image synthesis with reasonable quality by LoRA fine-tuning.
@@ 6.4 Application
One promising and attractive property of DMs is that they can be used in multiple downstream tasks. In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of DMs. Since the relevant applications are mostly developed for SD, we mainly show the results obtained by YOSO-LoRA.
Image-to-image. As shown in Fig. 6, our YOSO-LoRA is capable of performing high-quality image-to-image editing [32] in one step.
Compatibility with ControlNet and Different Base Models. As shown in Fig. 6, our YOSO-LoRA is compatible with ControlNet [64], following the condition well. Our YOSO-LoRA is also compatible with different base models (e.g., dreamshaper and realistic vision) fine-tuned from SD 1.5, preserving their style well. We find that when applying YOSO-LoRA to new base models, it fails to produce reasonable samples in one step. This may be due to the capability of LoRA and the distribution shift of training data.
| Model | NFE | HPS | AeS |
| SD 2.1-Turbo [44] | 1 | 27.06 | 5.83 |
| Realistic-vision | 50 | 29.76 | 6.35 |
| w/ YOSO-LoRA | 4 | 29.50 | 6.21 |
| w/ YOSO-LoRA-v2 | 1 | 27.13 | 6.12 |
| w/ YOSO-LoRA-v2 | 2 | 29.41 | 6.33 |
| Dreamshaper | 1 | 26.64 | 6.08 |
| w/ YOSO-LoRA-v2 |
Enabling Data Flywheel. YOSO-LoRA trained on the data generated by SDXL-Turbo, fails to generate reasonable samples in one step when compatible with new base models. In light of the fact that combining YOSO-LoRA with new base models can generate high-quality samples in 4 steps with minimal performance loss as well as preserved style, we propose using YOSO-generated data for training YOSO-LoRA based on existing community models (e.g., Realistic-vision). We call the trained model YOSO-LoRA-v2. As shown in Tab. 5, YOSO-LoRA-v2 can even outperform the full fine-tuned SD-2.1-Turbo in one step. This shows that the YOSO can be trained on slightly shifted data first, and then applied to generate high-quality un-shifted data to train a better YOSO. Moreover, the YOSO-LoRA-v2 can be compatible with some unseen base models (e.g., Dreamshaper) for one-step generation (see Figs. 6 and 5). This shows that combining YOSO-LoRA with community models can generate unshifted training data, enabling the YOSO-LoRA-v2 to focus on learning the acceleration with better compatibility.
@@ 7 Conclusion
Appendix A Broader Impacts
This work presents YOSO, a method that accelerates multi-step large-scale text-to-image diffusion models into a one-step generator. On the positive side, while this is academic research, we believe the proposed YOSO can be widely applied in industry, and its high efficiency can lead to energy savings and environmental benefits. However, when these rapid generation models are manipulated by malicious actors, they can also simplify and accelerate the creation of harmful information. Although our work focuses on scientific research, we will take actions to reduce the harmful information, such as filtering out harmful content in the dataset.
Appendix B Limitation
Our model, like most text-to-image diffusion models, may exhibit shortcomings in terms of fairness, as well as in handling specific details and accurately controlling the number of targets. We plan to explore these unresolved issues in the generation field in our future work, in order to enhance the model’s capabilities in text generation, fairness, detail control, and quantitative control.
Appendix C Safeguards
The YOSO-SD and YOSO-PixArt- are trained on the JourneyDB dataset [35], which has undergone rigorous human and machine-based filtering to ensure that there are no harmful or violent images in the dataset. YOSO-LoRA and YOSO-LoRA-V2, on the other hand, were trained using data generated by SDXL-turbo and YOSO-LoRA, respectively. To reduce the risk associated with the models, we utilized the built-in safety checker from diffusers [51] to filter out harmful data.
Appendix D Training Algorithm
We present the algorithm for training YOSO from scratch in Algorithm 1.
Appendix E Additional Related Work
Appendix F The Effect of Quick-Adaption
We present a qualitative comparison between YOSO w/o quick adaption and YOSO w/ quick adaption in Fig. 7. Note that the YOSO and variants here are trained with only 5k iterations. And for YOSO w/ quick adaption, we incorporate the 1k iterations used for quick adaptation into the total 5k training iterations to ensure a fair comparison. As shown in Fig. 7, YOSO w/ quick adaption demonstrates significantly better image quality compared to YOSO w/o quick adaption. The generated images of YOSO w/o quick adaption exhibit over-smoothing and severe artifacts. This indicates that the convergence of direct transition to v-prediction and zero terminal SNR is relatively slower. Additionally, it is worth highlighting that although there are only 5k training iterations, the one-step generation quality of YOSO is already reasonable, which demonstrates the effectiveness and quick convergence of the proposed quick adaption stage and YOSO.
Appendix G Experiment Setting Details
G.1 Unconditional Generation Experiments
For the generator, we use the Adam optimizer with = 0.9 and = 0.999; for the discriminator, we use the Adam optimizer with = 0. and = 0.999. We adopt a constant learning rate of 2e-4 for both the discriminator and the generator. We apply EMA with a coefficient of 0.9999 for the generator. We let the and .
G.2 Text-To-Image Experiments
For the generator, we use the AdamW optimizer with = 0.9 and = 0.999; for the discriminator, we use the AdamW optimizer with = 0. and = 0.999. We adopt a constant learning rate of 2e-5 for both the discriminator and the generator. We apply gradient norm clipping with a value of 1.0 for the generator only. We use batch size 256. For full fine-tuning, we apply EMA with a coefficient of 0.9999 for the generator. For LoRA fine-tuning, we apply EMA with a coefficient of 0.999 for the generator. Generally, the training is done within 30k iterations for full fine-tuning, while within 5k iterations for LoRA fine-tuning. We let the and .